A school has 18 classes with 35 students in each class. In order to reduce class size to 30, how many new classes must be formed?

Answer by stanbon(75887)

A school has 18 classes with 35 students in each class. In order to reduce class size to 30, how many new classes must be formed?
----------------
Total # of students = 18*35 = 630
------------------------------------
# of classes needed if 30 in a class: 630/30 = 21
--------------------------------------
# of new classes needed: 3
=================================
Cheers,
Stan H.